Rally Synthesis Inc. manufactures and sells 80 bottles per day. Fixed costs are $23,000 and the variable costs for manufacturing 80 bottles are $40,000.

Each bottle is sold for $1,500. How would the daily profit be affected if the daily volume of sales drop by 20​%?

A. profits are reduced $24,000

B. profits are reduced by $ $8,000

C.profits are reduced by $ $41,000

D.profits are reduced by $ 16,000
Answer #1

Given for Rally Synthesis Inc.

Fixed cost = $23000

Variable cost = $40000

So, variable cost per bottle = 40000/80 = $500 per bottle

Selling price = $1500 per bottle

So, contribution margin = 1500-500 = $1000 per bottle

Profit when 80 bottles are sold in a day = 80*1500 - 23000 = $97000

When sales drop by 20%, new sales = 80*0.8 = 64 bottles per day

So profit = 64*1500 - 23000 = $73000

So profits are reduced by $97000 - 73000 = $24000

So option A is correct.